I need help with these type of equations please,
-9h=-3(-3)

Answer by KnightOwlTutor(293)

Whenever you multiply a negative with another negative number the product is always positive. (-)(-)=(+)
multiply (-)(+)=(-)
-9h=-3(-3)therefore-9h=(+)9
h=9/(-)9=-1